OpenStudy (anonymous):

find the slope of a tangent line to the parametric curve x(t) = 4t - 4sint, y(t) = 4-4cost when t=pi/4. Use dy/dx = (dy/dt)/(dx/dt)

OpenStudy (anonymous):

I got 1 for an answer, if someone could do and check I would appreciate!

OpenStudy (anonymous):

x(t) = 4t - 4sint, y(t) = 4-4cost when t=pi/4. Use dy/dx = (dy/dt)/(dx/dt) Dx(t) = 4 - 4cost Dy(t)=4sin t now dy/dx=(dy/dt)/(dx/dt) =(4sin t)/(4 - 4cost) using t=pi/4 =1+sqrt2

OpenStudy (anonymous):

cool can you show work from last step to second to last step

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

now dy/dx=(dy/dt)/(dx/dt) =(4sin t)/(4 - 4cost) using t=pi/4 =(4sin (pi/4))/(4 - 4cos(pi/4)) =4(sqrt2/2)/4 - 4(sqrt2/2))) =2sqrt2/(4-2sqrt2) mult both sides of the eq by (4+2sqrt2) 2sqrt2 (4+2sqrt2) =---------------------- (4-2sqrt2) (4-2sqrt2) 8sqrt2 +4(2) =------------ 16-4(2) 8sqrt2+8 =---------- dividing by 8 we get 8 =sqrt2+1 ans

OpenStudy (anonymous):

any question ivn?

OpenStudy (anonymous):

on mult both sides by (4+2sqrt2) youll noticed i made it negative sign,its positive sign,lol

OpenStudy (anonymous):

ok nice

OpenStudy (anonymous):

ok have fun now, good luck

OpenStudy (anonymous):

i have another question reguarding parametrics im about to post oif you could help with that